OpenStudy (tumblewolf):

Will Fan and medal Simple linear algebra!! Write the standard form of the line that passes through the given points. 1. (1, 5) and (-2, 3) 2. (-2, 4) and parallel to x-2y=6 3. a slope of 1/2 and a y intercept of 3 (For this one it's set up as y=1/2x+3 but I don't know how to convert it over to standard) 4. (-1, -3) (2,1)

OpenStudy (bibby):

do you know what point slope form looks like?

OpenStudy (tumblewolf):

y=m(x+b)

OpenStudy (tumblewolf):

@bibby

OpenStudy (bibby):

\(y-y_1=m(x-x_1)\) where \((x_1,y_1)\) is a point the line goes through and m is the slope